OpenStudy (anonymous):

lim x->0^+ (tan6x)^x

OpenStudy (anonymous):

the question is not very clear?

OpenStudy (anonymous):

you may take log of function and the use hopital rule

OpenStudy (anonymous):

\[y=(\tan 6x)^x\]\[\ln y=x \ln 6x\]\[\ln y= \frac{\ln 6x}{(1/x)}\]now take the limit of RHS when x->0+

OpenStudy (anonymous):

for checking answer is:\[1^-\]